Week of 170807

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:

  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Kate (chair, DB), Andrew (LHCb), Jiri (ATLAS), Nils (computing), Herve (storage), Roberto (storage), Marian (networks)
  • remote: Luca (CNAF), Chihsun (ASGC), Andrew (NL-T1), Sang Un (KISTI), Dmytro (NDGF), Daniele (CMS), Kyle (OSG), Gareth (RAL), Dave (FNAL)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Activities:
      • Grid production continuing with number of used slots mostly around 300k, short dip to 120k slots because of lack of available tasks
    • Issues:
      • Even when number of overlay jobs was limited to 1200, we observed issues with frontier servers in RAL and IN2P3. The first solved by restart, the latter was just a monitoring issue. Now we have set the limit to 800 overlay jobs, no problems with frontier servers observed anymore.
      • ATLASEOS various problems, performance issues solved by adding more servers last week and small changes in configuration

  • CMS reports ( raw view) -
    • excellent LHC week-end for LHC. Saturday, a nice long fill (#6050, ~15hrs). Sunday, fill #6052 good too
    • Machine availability was 95%, 65% in Stable Beams + 5% in adjust, 1.7 fb-1 since Friday morning for a total of 13.6 fb-1
    • CMS is taking data just fine
    • prompt reco paused at the T0 because of a missing prompt-calib loop for a run (being worked on, and of limited interest to this meeting)
    • Run2017 MC ramping up in next days. Already visible activity from Phase II Upgrade samples
    • T0: GGUS:129937: 1-file loss on EOS, causing the corresponding Run2017C express job failing
    • T1: GGUS:129944: load on PIC, keeping up but heavy (will be discussed at the facilities and ops meetings later today)
    • T1: GGUS:129954: transfer quality to T1_FR_CCIN2P3(_disk) indicates failure of all transfers

  • ALICE -
    • Apologies: ALICE operations will not attend today
    • Very high activity
      • New concurrent job records were reached, up to 155k on Thu
    • CERN:
      • New HTCondor CE hosts ce509 ... ce514 are OK so far, thanks!
      • ALICE jobs by default no longer request any output sandbox
        • thus reducing I/O by a lot

  • LHCb reports ( raw view) -
    • Activity
      • Monte Carlo simulation, data processing and user analysis
    • Site Issues
      • T0:
        • Key VO Box (lbvobox103) unavailable (lost?) due to hypervisor problem. (GGUS:129942). No GGUS (or Service Now) updates since yesterday morning. Having to recreate services on other VO Boxes. Nils will check the missing updates.
      • T1:
        • NTR

Sites / Services round table:

  • ASGC: NTR
  • BNL: NTR
  • CNAF: NTR
  • EGI: NC
  • FNAL: NTR
  • IN2P3: NC
  • JINR: NC
  • KISTI: NTR
  • KIT: NC
  • NDGF: NTR
  • NL-T1: NTR
  • NRC-KI: NC
  • OSG: NTR
  • PIC: NC
  • RAL: Problems with the Atlas Castor SRMs in the early morning of Saturday 5th August. The disk pools for AtlasDataDisk and AtlasScratchDisk in Castor have been merged.
  • TRIUMF:NTR

  • CERN computing services: NTR.
  • CERN storage services:
    • EOSATLAS and EOSCMS: Bug that could cause file loss under special failure circumstances even though EOS acknowledges the write to the client. Fix ready, and deploying on both EOSATLAS and EOSCMS (will take days to apply to all nodes in a non disruptive manner)
CMS will take more time than ATLAS as the number of writes in ATLAS is lower due to Thursday EOS issue.
  • CERN databases: NTR
  • GGUS: NTR
  • Monitoring:
    • Draft reports for the July 2017 availability sent around
  • MW Officer: NC
  • Networks:
    • CBPF reported transfer problems to CNAF, PIC and IN2P3 after RedCLARA decommissioned Sao Paulo - Madrid direct circuit - looks like MTU step down issue
    • RO-02-NIPNE to multiple sites (GGUS:128489) - after investigating it doesn't appear to be a network issue, but additional checks are needed by site
  • Security: NC

AOB:

Edit | Attach | Watch | Print version | History: r17 < r16 < r15 < r14 < r13 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r17 - 2017-08-07 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback